OpenStudy (anonymous):

What is the value of x? A. 4 B. 5 C. 8 D. 10

hartnn (hartnn):

All 3 sides are given equal so equate them \(5x-9=3x+1\) can you try to find x from here ?

OpenStudy (anonymous):

so I subtract 5x-9=-4? and then 3x+1=4x?

hartnn (hartnn):

you cannot combine 5x and -9 they are not like terms!

OpenStudy (anonymous):

oh right like terms

hartnn (hartnn):

\(5x-9 = 3x+1\) add 9 on both sides \(5x -9+9 = 3x+1+9 \\ 5x =3x+10 \) did u get this ?

OpenStudy (anonymous):

yea i am getting the idea

hartnn (hartnn):

now try to subtract 3x from both sides what do u get ?

OpenStudy (anonymous):

hmmm... I would said C.? I think

hartnn (hartnn):

no...see the steps \(5x =3x +10 \\ 5x -3x = 3x-3x+10 \\ 2x = 10 \\ x=5\) see if you get them, ask if any doubts

OpenStudy (anonymous):

oh Thank you sooooo much!!

hartnn (hartnn):

welcome ^_^